# Algebraic and geometric properties of homeomorphism groups of ordinals

### v1

@@ 4.2 Properties of $\overline{\operatorname{F}}_{0,d}$
We stated the fragmentation lemma for general $\alpha$, but it is particularly useful for $\alpha=1$, as in this case $\operatorname{F}_{0,d}$ is the subgroup of finitely supported homeomorphisms.
@@
Let $G$ denote the subgroup of $\overline{\operatorname{F}}_{\alpha,d}$ consisting of elements supported in $U$. We can then identify $G$ with a subgroup of $\operatorname{Homeo}(U)$. By 2.18, $U$ is homeomorphic to $X_{\alpha,1}$, and hence as $G$ contains an element inducing an infinite permutation on the set $N_{\alpha,d}$, Theorem 3.12 implies $G$ is isomorphic to $\operatorname{Homeo}(\omega^{\alpha+1})$. ∎
4.5 and 4.6 imply that $\overline{\operatorname{F}}_{\alpha,d}/\operatorname{F}_{\alpha,d}$ can be realized as quotient of $\operatorname{Homeo}(\omega^{\alpha+1})^{d}$, and hence this quotient is strongly distorted and uniformly perfect. In the case of uniform perfectness, the difficulty is then understanding whether we can express the elements of $\operatorname{F}_{\alpha,d}$ as products of commutators of elements in $\overline{\operatorname{F}}_{\alpha,d}$. Unfortunately, we cannot answer this question in general, but we can do so for $\alpha=0$.
@@
We now turn to classifying the normal generators of $\overline{\operatorname{F}}_{0,d}$. Given $Z\subset M_{0,d}$, let $\operatorname{F}_{Z}$ denote the subgroup of $\overline{\operatorname{F}}_{0,d}$ consisting of homeomorphisms that restrict to the identity in an open neighborhood of each point of $Z$. Given a subgroup $\Gamma$ of $\overline{\operatorname{F}}_{0,d}$, we let $Z_{\Gamma}$ be the subset of $M_{0,d}$ consisting of the $\mu\in M_{0,d}$ such that for each $g\in\Gamma$ there is an open neighborhood $U$ of $\mu$ that is fixed pointwise by $g$, that is, $Z_{\Gamma}$ is the largest subset of $M_{0,d}$ satisftying $\operatorname{F}_{Z_{\Gamma}}<\Gamma$. For $f\in\overline{\operatorname{F}}_{0,d}$, we write $Z_{f}$ to denote $Z_{\langle f\rangle}$.
@@
Let $d\in\mathbb{N}\smallsetminus\{1\}$, and let $\Gamma$ be a normal subgroup of $\overline{\operatorname{F}}_{0,d}$. If $\Gamma$ is not contained in $\operatorname{F}_{0,d}$, then $\operatorname{F}_{Z_{\Gamma}}=\Gamma$. Moreover, (1) there exists $\gamma\in\Gamma$ such that $Z_{\gamma}=Z_{\Gamma}$, and (2) if $\gamma\in\Gamma$ such that $Z_{\gamma}=Z_{\Gamma}$, then $\gamma$ uniformly normally generates $\Gamma$.
@@
By assumption, $\Gamma$ is a normal subgroup of $\overline{\operatorname{F}}$ not contained in $\operatorname{F}$, and therefore, $Z_{\Gamma}\neq M$. As the elements of $N_{0,d}$ are isolated, it follows that $\Gamma<F_{Z_{\Gamma}}$, so we need only show that $F_{Z_{\Gamma}}<\Gamma$.
First, we claim that if $\mu\in M\smallsetminus Z_{\Gamma}$, then there exists a clopen neighborhood $U_{\mu}$ of $\mu$ and an element $g_{\mu}$ in $\Gamma$ supported in $U_{\mu}$ such that $U_{\mu}\cap M=\{\mu\}$ and such that $g_{\mu}$ induces an infinite permutation on $N\cap U_{\mu}$. Fix $\mu\in M\smallsetminus Z_{\Gamma}$, and choose $g\in\Gamma$ such that $g$ induces a nontrivial permutation on $U\cap N$ for every open neighborhood $U$ of $\mu$. By continuity, we can choose pairwise-disjoint clopen neighborhoods $U_{1},\ldots,U_{d}$ of $\mu_{1},\ldots,\mu_{d}$, respectively, such that $g(U_{i})\cap U_{j}=\varnothing$ whenever $j\neq k$. Applying 4.5 to $g$ allows us to write $g=g_{0}\circ g_{1}\circ\cdots\circ g_{d}$ with $g_{0}\in\operatorname{F}$ and with $g_{i}\in\overline{\operatorname{F}}$ supported in $U_{i}$.
Let $k\in\{1,\ldots,d\}$ be such that $\mu=\mu_{k}$. By 3.11, there exists an infinite subset $A$ of $N\cap U_{k}$ such that $g_{k}(A)\cap A=\varnothing$. By shrinking $A$ we may assume that $A$ is disjoint from the support of $g_{0}$. Let $\tau$ be the involution such that $\tau|_{A}=g_{k}|_{A}$, $\tau|_{g_{k}(A)}=g_{k}^{-1}|_{g_{k}(A)}$, and such that $\tau$ is the identity on the complement of $A\cup g_{k}(A)$. It follows that $[\tau,g_{0}g_{k}]$ induces an infinite permutation on $U_{k}\cap N$, and
@@
$=[\tau,g_{0}g_{k}]$
The first equality implies that $[\tau,g]\in\Gamma$, and the second equality is deduced from the fact that $\tau$ commutes with each $g_{i}$ for $i>0$ and $i\neq k$. As $\tau$ is supported in $U_{k}$, it follows from the second equality that $[\tau,g]$ restricts to the identity on the complement of $g_{0}(U_{k})$, i.e., $[\tau,g]$ is supported in $g_{0}(U_{k})$. Setting $g_{\mu}=[\tau,g]$ and $U_{\mu}=g_{0}(U_{k})$, establishes the claim.
Let $G_{\mu}$ denote the subgroup of $\overline{\operatorname{F}}$ consisting of elements supported in $U_{\mu}$. 4.6 implies that $G_{\mu}$ is isomorphic to $\operatorname{Homeo}(\omega^{\alpha+1})$. In particular, 3.12 implies that $g_{\mu}$ uniformly normally generates $G_{\mu}$. Moreover, as $g_{\mu}\in\Gamma$ and $\Gamma$ is normal, we have that $G_{\mu}<\Gamma$.
Now, fix $f\in\operatorname{F}_{Z_{\Gamma}}$. Again applying 4.5, we can write $f=f_{0}\circ\prod_{\mu\in M\smallsetminus Z_{\Gamma}}f_{\mu}$ with $f_{0}\in\operatorname{F}$ and each $f_{\mu}\in\overline{\operatorname{F}}$ supported in $U_{\mu}$. It follows that there exist $w\in\overline{\operatorname{F}}_{0,d}$ such that $f\circ w=f_{0}\in\operatorname{F}$ and such that $w$ can be expressed as a product of $12|M\smallsetminus Z_{\Gamma}|$ elements from the set $\{g_{\mu},g_{\mu}^{-1}:\mu\in M\smallsetminus Z_{\Gamma}\}$ and their conjugates (the number twelve comes from the proof of 3.12). In particular, $w\in\Gamma$. It is left to check that $f_{0}\in\Gamma$. As $f_{0}\in\operatorname{F}_{0,d}$ has finite support, there is a conjugate of $f_{0}$ supported in $U_{\mu_{1}}$. 3.9 implies that $f_{0}$ can be written as a product of four conjugates of $g_{\mu_{1}}$ and $g_{\mu_{1}}^{-1}$. Therefore, $F_{Z_{\Gamma}}<\Gamma$, as desired.
@@ 4.3 Properties of $\operatorname{PH}_{0,d}$ and $\operatorname{H}_{0,d}$
Let $\alpha$ be an ordinal, and let $d\in\mathbb{N}$. The minimal cardinality for a normal generating set of $\operatorname{PH}_{0,d}$ is $d-1$ and is two for $\operatorname{H}_{0,d}$.
@@
Let us first consider $\operatorname{PH}_{0,d}$. First note that, by 4.10, the cardinality of any normal generating set for $\operatorname{PH}_{0,d}$ must be at least $d-1$, as it projects to a generating set for $\mathbb{Z}^{d-1}$. So, we need only find a normal generating set of cardinality of $d-1$; to this end, let $S=\{s_{1},\ldots,s_{d-1}\}$, where the $s_{i}$ are the shift homeomorphisms constructed in the proof of 4.1. We can then label the elements of $N$ as $\{(n,i):n\in\mathbb{N},1\leq i\leq d-1\}$ in such a way that $s_{i}(n,i)=(n+1,i)$ and $s_{i}(n,j)=(n,j)$ whenever $j\neq i$. Let $s=s_{1}\circ s_{2}\circ\cdots\circ s_{d-1}$. Choose $t\in\operatorname{PH}_{0,d}$ such that $t(2n,i)=(2n+1,i)$ and $t(2n+1,i)=(2n,i)$. Then $s^{t}(2n,i)=(2n+3,i)$ and $s^{t}(2n+1,i)=(2n,i)$. In particular,
$[t,s](2n+1,i)=(2n+3,i)\text{ and }[t,s](2n,i)=(2n-2,i)$
It follows that $[t,s]\in\overline{\operatorname{F}}_{0,d}$ and $Z_{[t,s]}=\varnothing$; therefore, $[t,s]$ normally generates $\overline{\operatorname{F}}_{0,d}$ by 4.9. By 4.4, $\operatorname{PH}_{0,d}$ is the semi-direct product of $\overline{\operatorname{F}}_{0,d}$ and the subgroup generated by $S$, and hence $S$ is a normal generating set for $\operatorname{PH}_{0,d}$.

### v2

@@ 4.2 Properties of $\overline{\operatorname{F}}_{0,d}$
We stated the fragmentation lemma for general $\alpha$, but it is particularly useful for $\alpha=0$, as in this case $\operatorname{F}_{0,d}$ is the subgroup of finitely supported homeomorphisms.
@@
Let $G$ denote the subgroup of $\overline{\operatorname{F}}_{\alpha,d}$ consisting of elements supported in $U$. We can then identify $G$ with a subgroup of $\operatorname{Homeo}(U)$. By 2.21, $U$ is homeomorphic to $X_{\alpha,1}$, and hence as $G$ contains an element inducing an infinite permutation on the set $N_{\alpha,d}$, Theorem 3.12 implies $G$ is isomorphic to $\operatorname{Homeo}(\omega^{\alpha+1})$. ∎
4.5 and 4.6 imply that $\overline{\operatorname{F}}_{\alpha,d}/\operatorname{F}_{\alpha,d}$ can be realized as a quotient of $\operatorname{Homeo}(\omega^{\alpha+1})^{d}$, and hence this quotient is strongly distorted and uniformly perfect. In the case of uniform perfectness, the difficulty is then understanding whether we can express the elements of $\operatorname{F}_{\alpha,d}$ as products of commutators of elements in $\overline{\operatorname{F}}_{\alpha,d}$. Unfortunately, we cannot answer this question in general, but we can do so for $\alpha=0$.
@@
We now turn to classifying the normal generators of $\overline{\operatorname{F}}_{0,d}$. Given $Z\subset M_{0,d}$, let $\operatorname{F}_{Z}$ denote the subgroup of $\overline{\operatorname{F}}_{0,d}$ consisting of homeomorphisms that restrict to the identity in an open neighborhood of each point of $Z$. Given a subgroup $\Gamma$ of $\overline{\operatorname{F}}_{0,d}$, we let $Z_{\Gamma}$ be the subset of $M_{0,d}$ consisting of the $\mu\in M_{0,d}$ such that for each $g\in\Gamma$ there is an open neighborhood $U$ of $\mu$ that is fixed pointwise by $g$, that is, $Z_{\Gamma}$ is the largest subset of $M_{0,d}$ satisfying $\operatorname{F}_{Z_{\Gamma}}<\Gamma$. For $f\in\overline{\operatorname{F}}_{0,d}$, we write $Z_{f}$ to denote $Z_{\langle f\rangle}$.
@@
Let $d\in\mathbb{N}\smallsetminus\{1\}$, and let $\Gamma$ be a normal subgroup of $\overline{\operatorname{F}}_{0,d}$. If $\Gamma$ is not contained in $\operatorname{F}_{0,d}$, then $\operatorname{F}_{Z_{\Gamma}}=\Gamma$. Moreover, (1) there exists $\gamma\in\Gamma$ such that $Z_{\gamma}=Z_{\Gamma}$, and (2) if $\gamma\in\Gamma$ is such that $Z_{\gamma}=Z_{\Gamma}$, then $\gamma$ uniformly normally generates $\Gamma$.
@@
By assumption, $\Gamma$ is a normal subgroup of $\overline{\operatorname{F}}$ not contained in $\operatorname{F}$, and therefore, $Z_{\Gamma}\neq M$. As the elements of $N=N_{0,d}$ are isolated, it follows that $\Gamma\leq F_{Z_{\Gamma}}$, so we need only show that $F_{Z_{\Gamma}}\leq\Gamma$.
First, we claim that if $\mu$ belongs to $M$ but not $Z_{\Gamma}$, then there exists a clopen neighborhood $U_{\mu}$ of $\mu$ and an element $g_{\mu}$ in $\Gamma$ supported in $U_{\mu}$ such that $U_{\mu}\cap M=\{\mu\}$ and such that $g_{\mu}$ induces an infinite permutation on $N\cap U_{\mu}$.
Indeed, fix $\mu\in M\smallsetminus Z_{\Gamma}$, and choose $g\in\Gamma$ such that $g$ induces a nontrivial permutation on $U\cap N$ for every open neighborhood $U$ of $\mu$. By continuity, we can choose pairwise-disjoint clopen neighborhoods $U_{1},\ldots,U_{d}$ of $\mu_{1},\ldots,\mu_{d}$, respectively, such that $g(U_{i})\cap U_{j}=\varnothing$ whenever $j\neq k$. Applying 4.5 to $g$ allows us to write $g=g_{0}\circ g_{1}\circ\cdots\circ g_{d}$ with $g_{0}\in\operatorname{F}$ and with $g_{i}\in\overline{\operatorname{F}}$ supported in $U_{i}$.
Let $k\in\{1,\ldots,d\}$ be such that $\mu=\mu_{k}$. By 3.11, there exists an infinite subset $A$ of $N\cap U_{k}$ such that $g_{k}(A)\cap A=\varnothing$. By shrinking $A$ we may assume that $A$ is disjoint from the support of $g_{0}$. Let $\tau$ be the involution such that $\tau|_{A}=g_{k}|_{A}$, $\tau|_{g_{k}(A)}=g_{k}^{-1}|_{g_{k}(A)}$, and such that $\tau$ is the identity on the complement of $A\cup g_{k}(A)$. It follows that $\tau,g_{0}g_{k}]$ induces an infinite permutation on $U_{k}\cap N$, and
@@
$=[\tau,g_{0}g_{k}].$
The first equality implies that $\tau,g]\in\Gamma$, and the second equality is deduced from the fact that $\tau$ commutes with each $g_{i}$ for $i>0$ and $i\neq k$. As $\tau$ is supported in $U_{k}$, it follows from the second equality that $\tau,g]$ restricts to the identity on the complement of $g_{0}(U_{k})$, i.e., $\tau,g]$ is supported in $g_{0}(U_{k})$. Setting $g_{\mu}=[\tau,g]$ and $U_{\mu}=g_{0}(U_{k})$, establishes the claim.
Now, let $G_{\mu}$ denote the subgroup of $\overline{\operatorname{F}}$ consisting of elements supported in $U_{\mu}$. 4.6 implies that $G_{\mu}$ is isomorphic to $\operatorname{Homeo}(\omega^{\alpha+1})$. In particular, 3.12 implies that $g_{\mu}$ uniformly normally generates $G_{\mu}$. Moreover, as $g_{\mu}$ belongs to $\Gamma$ and $\Gamma$ is normal, we have that $G_{\mu}$ is a subgroup of $\Gamma$.
Next, fix $f\in\operatorname{F}_{Z_{\Gamma}}$. Again applying 4.5, we can write $f=f_{0}\circ\prod_{\mu\in M\smallsetminus Z_{\Gamma}}f_{\mu}$ with $f_{0}\in\operatorname{F}$ and each $f_{\mu}\in\overline{\operatorname{F}}$ supported in $U_{\mu}$. It follows that there exists $w\in\overline{\operatorname{F}}_{0,d}$ such that $f\circ w=f_{0}\in\operatorname{F}$ and such that $w$ can be expressed as a product of $12|M\smallsetminus Z_{\Gamma}|$ elements from the set $\{g_{\mu},g_{\mu}^{-1}:\mu\in M\smallsetminus Z_{\Gamma}\}$ and their conjugates (the number twelve comes from the proof of 3.12). In particular, $w$ belongs to $\Gamma$. It is left to check that $f_{0}$ is an element of $\Gamma$. As $f_{0}\in\operatorname{F}_{0,d}$ has finite support, there is a conjugate of $f_{0}$ supported in $U_{\mu_{1}}$. 3.9 implies that $f_{0}$ can be written as a product of four conjugates of $g_{\mu_{1}}$ and $g_{\mu_{1}}^{-1}$. Therefore, $F_{Z_{\Gamma}}$ is a subgroup of $\Gamma$, as desired.
@@ 4.3 Properties of $\operatorname{PH}_{0,d}$ and $\operatorname{H}_{0,d}$
Let $\alpha$ be an ordinal, and let $d\in\mathbb{N}$. The minimal cardinality for a normal generating set of $\operatorname{PH}_{0,d}$ is d1d-1 and is two for $\operatorname{H}_{0,d}$.
@@
Let us first consider $\operatorname{PH}_{0,d}$. First note that, by 4.10, the cardinality of any normal generating set for $\operatorname{PH}_{0,d}$ must be at least d1d-1, as it projects to a generating set for $\mathbb{Z}^{d-1}$. So, we need only find a normal generating set of cardinality of d1d-1; to this end, let $S=\{s_{1},\ldots,s_{d-1}\}$, where the $s_{i}$ are the shift homeomorphisms constructed in the proof of 4.1. We can then label the elements of $N$ as $\{(n,i):n\in\mathbb{N},1\leq i\leq d-1\}$ in such a way that $s_{i}(n,i)=(n+1,i)$ and $s_{i}(n,j)=(n,j)$ whenever $j\neq i$. Let $s=s_{1}\circ s_{2}\circ\cdots\circ s_{d-1}$. Choose $t\in\operatorname{PH}_{0,d}$ such that $t(2n,i)=(2n+1,i)$ and $t(2n+1,i)=(2n,i)$. Then $s^{t}(2n,i)=(2n+3,i)$ and $s^{t}(2n+1,i)=(2n,i)$. In particular,
$t,s](2n+1,i)=(2n+3,i)\text{ and }[t,s](2n,i)=(2n-2,i)$
It follows that $t,s]\in\overline{\operatorname{F}}_{0,d}$ and $Z_{[t,s]}=\varnothing$; therefore, $[t,s]$ normally generates $\overline{\operatorname{F}}_{0,d}$ by 4.9. By 4.4, $\operatorname{PH}_{0,d}$ is the semi-direct product of $\overline{\operatorname{F}}_{0,d}$ and the subgroup generated by $S$, and hence $S$ is a normal generating set for $\operatorname{PH}_{0,d}$.